Grade 12th passMechanics

two bodies of unequal masses are thrown upwards with same velocity.if air exerts equal force on both bodies,will they reach same height or not?explain.?

Rakesh Hotker
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
let force exert by air is F on both the bodies

do retardation = (F+mg)/m = g + F/m

clearly retardation on lighter body will  be more .so it will reach less hight as compair to heavier body.